evaluate ∫ (3x sqrt(x) + 4 sqrt(x) + 5) dx

evaluate
∫ (3x sqrt(x) + 4 sqrt(x) + 5) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (3x\sqrt{x} + 4\sqrt{x} + 5)dx
I = \int (3x^{3/2} + 4\sqrt{x} + 5)dx
I = \int 3x^{3/2}dx + \int 4\sqrt{x}dx + \int 5dx
I = \frac{3x^{5/2}}{5/2} + \frac{4x^{3/2}}{3/2} + 5x + constant
I = \frac{6x^{5/2}}{5} + \frac{8x^{3/2}}{3} + 5x + constant